3 digit number divisible by 3 and two digits repeated.....???

Grade:11

2 Answers

Aman Bansal
11 years ago

DearShubham,

For the number to be divisible by 4, the number formed by the last two digits must be divisible by 4. With the digits 1 through 6, you can use the 3 last digits 24 and 6, and each of these can be combined with half of the digits to form a 2-digit number divisible by 4, so there are 9combinations for the last two digits. The other two digits can be chosen arbitrarily, so there are 36combinations for those, for a total of 9\cdot 36=324 different numbers

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Peetala Kishore
11 years ago

Hey

suppose 114

to test this = 1+1+4 =6 .

six is divisable by 3 so

the given number also divisable by 3.

there are 66 possiblities are there for your question.

they are = {114,117,225,228,330,336,339,441,448,552,558,660,669,771,774,882,885,990,993,996} there is also possibloe of repetitions  other are 300,600,900.